OpenStudy (anonymous):

Henry grabs one shoe at a time at random without replacement from a box of 14 shoes in his closet. The box contains 6 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be brown? Round to the nearest percent if necessary.

OpenStudy (cwtan):

6/14 *5/13

Parth (parthkohli):

P(Brown shoe) = 6/14 = 3/7 If you want it both times, then multiply the probabilities. 3/7 * 3/7

Parth (parthkohli):

Oops

Parth (parthkohli):

3/7 * 5/13 because no replacement is done